find x
when the hypotenuse is x, the long leg is (x-2), the short leg is 8

Answer by stanbon(75887)

find x
when the hypotenuse is x, the long leg is (x-2), the short leg is 8
-------
Equation:
8^2 + (x-2)^2 = x^2
--------------
64 + x^2-4x+4 = x^2
--------------
4x = 60
x = 15
-----------
Cheers,
Stan H.
===========